# Dynamic Programming: From Local Optimality to Global Optimality

### v2

@@ Abstract.
In the theory of dynamic programming, an optimal policy is a policy whose lifetime value dominates that of all other policies from every possible initial condition in the state space. This raises a natural question: when does optimality from a single state imply optimality from every state? We show that, in a general setting, irreducibility of the transition kernel is sufficient for this property. Our results have important implications for modern policy-based algorithms used to solve large-scale dynamic programs in reinforcement learning and other fields.
@@ 1. Introduction
Dynamic programming (DP) is a major branch of optimization theory, with applications ranging from fine tuning of large language models to DNA sequencing, space exploration, and air traffic control. Dynamic programs that include uncertainty are often called Markov decision processes (MDPs) and the theory of such processes has been extensively developed (see, e.g., Bäuerle and Rieder (2011), Hernández-Lerma and Lasserre (2012), Bertsekas (2012), or Bertsekas (2022)). Much of the recent surge in interest in MDPs has been fueled by artificial intelligence and reinforcement learning (see, e.g., Bertsekas (2021) or Kochenderfer et al. (2022)).
In recent years, researchers solving large-scale MDPs have moved away from value-based methods and towards policy-based methods, one example of which is policy gradient ascent (see, e.g., Murphy (2024), Sutton et al. (1999), Lan et al. (2023), Kumar et al. (2023), or Friedl et al. (2023)). These policy-based methods seek to maximize a real-valued objective, such as
@@
Despite these successes, there is one significant disadvantage of policy gradient methods: unlike some traditional DP algorithms, these methods are not guaranteed to find an optimal policy. To understand the issues at hand, recall that an optimal policy is a feasible policy $\sigma$ such that $v_{\sigma}(x)=\max_{s\in\Sigma}v_{s}(x)$ for every $x\in\mathsf{X}$; that is, a policy $\sigma$ such that following this policy in every period leads to maximum lifetime value from every initial state $x$. Even if one attains a global maximum in (1), with maximization over all $\sigma\in\Sigma$, there is no guarantee that the resulting policy will be an optimal policy.
The prevailing view is that obtaining an approximately optimal policy depends heavily on the initial distribution $\rho$ that is used to construct the objective function (1). For example, in an important study of policy gradient methods, Bhandari and Russo (2024) state that “Policy gradient methods have poor convergence properties if applied without an exploratory initial distribution” (Bhandari and Russo, 2024, p. 1910). Here “exploratory” means that $\rho$ in (1) should have a large support.
One disadvantage of choosing $\rho$ with large support is that this exploration can be computationally expensive. This motivates the following question: When is it permissible to choose $\rho$ with small support? At the extreme, when can $\rho$ be concentrated at a single point $x$, so that the maximization criterion $m(\sigma)$ from (1) is just $v_{\sigma}(x)$, and yet maximization of this criterion over all $\sigma\in\Sigma$ yields a (globally) optimal policy? In other words, when does optimality at a single state imply optimality at every state? We show that, for standard MDPs on general state spaces, irreducibility of the Markov dynamics generated by $\sigma$ is sufficient for this property. Specifically, if a policy is optimal at a single state and has an irreducible transition kernel, then this optimality propagates throughout the entire state space, making the policy globally optimal. Similarly, if irreducibility holds and there exists a distribution $\rho$ such that $\int v_{\sigma}\mathop{}\!\mathrm{d}\rho=\max_{s\in\Sigma}\int v_{s}\mathop{}\!\mathrm{d}\rho$, then $\sigma$ is an optimal policy. In addition, we show that two weak forms of irreducibility are sufficient for this result when $\sigma$ is also continuous.
@@
Papers examining theoretical properties of policy gradient methods have some connection to our work. For example Bhandari and Russo (2024) examine when policy gradient ascent (actually decent) yields a globally optimal policy. One the one hand, Bhandari and Russo (2024) directly examine the policy gradient algorithm and find sharp new results for several important cases. On the other hand, to step from local to global optimality, they restrict the classes of MDPs under consideration and assume a large support for $\rho$. Here we avoid large support restrictions on $\rho$ (since imposing such restrictions directly imposes a connection from local to global optimality—see Lemma 3.2 and the surrounding discussion).
Other papers that examine theoretical properties of gradient policy methods include Khodadadian et al. (2021), Agarwal et al. (2021), and Xiao (2022). However, in these papers, the focus is on proving the convergence of $\int v_{\sigma}\mathop{}\!\mathrm{d}\rho$ to its maximal value, rather than obtaining global convergence from local convergence (as we do here). At the same time, these papers provide valuable rates of convergence for specific algorithms, which we do not discuss. A related line of research focuses on average-optimal policies in finite-state MDPs by leveraging specific state space structures under some policies. This includes exploring unichain, multichain, communicating, and weakly communicating MDPs to study algorithmic convergence (Bartlett and Tewari, 2009; Puterman, 2014). Our results also demonstrate that, in finite-state MDPs, optimality can extend from a single state to all accessible states. Thus, our results are applicable to various classes of MDPs in this line of research and support the development of more efficient algorithms.
The paper is structured as follows. Section 2 provides background on MDPs. Section 3 states our main results in a general setting. Section 4 examines how the irreducibility condition from Section 3 can be weakened while still obtaining some transmission of optimality across states. Section 5 illustrates our theoretical results in the context of a benchmark optimal savings problem. Section 6 outlines avenues for future work.
@@ 2.1. Preliminaries
A transition kernel on $\mathsf{X}$ is a function $P$ from $\mathsf{X}\times\mathscr{B}$ to $[0,1]$ such that $x\mapsto P(x,B)$ is Borel measurable for all $B\in\mathscr{B}$ and $B\mapsto P(x,B)$ is a Borel probability measure for all $x\in\mathsf{X}$. To any such transition kernel $P$ we associate a bounded linear operator on $b\mathsf{X}$, often referred to as its Markov operator and also denoted by $P$, via
$f\mapsto Pf,\qquad(Pf)(x)=\int f(x^{\prime})P(x,\mathop{}\!\mathrm{d}x^{\prime}).$ (2)
In what follows, $(Pf)(x)$ will be understood as the expectation of $f(X_{t+1})$ given that $X_{t}=x$ and $X_{t+1}$ is drawn from $P(x,\mathop{}\!\mathrm{d}x^{\prime})$.
As usual, the positive cone of $b\mathsf{X}$, denoted here by $b\mathsf{X}_{+}$, is the set of all nonnegative functions in $v\in b\mathsf{X}$. Let $b\mathsf{X}^{\prime}$ be the dual space of $b\mathsf{X}$ and let $b\mathsf{X}^{\prime}_{+}$ to be the positive cone of $b\mathsf{X}^{\prime}$. The set $b\mathsf{X}^{\prime}_{+}$ contains, among other objects, the set $\mathcal{D}(\mathsf{X})$ of Borel probability measures on $\mathsf{X}$. For simplicity, elements of $\mathcal{D}(\mathsf{X})$ are referred to as distributions. For $\rho\in\mathcal{D}(\mathsf{X})$ and $f\in b\mathsf{X}$ we set
@@
For each $x\in\mathsf{X}$, the point evaluation functional generated by $x$ is the map $\delta_{x}$ that sends each $w\in b\mathsf{X}$ into $w(x)\in\mathbbm{R}$. Below it will be convenient for us to write this in dual notation, so that $\left\langle{w,\delta_{x}}\right\rangle=w(x)$ for every $w\in b\mathsf{X}$. We will make use of the following lemma.
@@
Regarding continuity, if $w_{n}\to w$ in $b\mathsf{X}$, then $w_{n}\to w$ pointwise on $\mathsf{X}$, so $\left\langle{w_{n},\delta_{x}}\right\rangle=w_{n}(x)\to w(x)=\left\langle{w,\delta_{x}}\right\rangle$. Regarding positivity, it suffices to show that $\left\langle{w,\delta_{x}}\right\rangle\geqslant 0$ whenever $w\geqslant 0$. This clearly holds, since $w\geqslant 0$ implies $w(x)=\left\langle{w,\delta_{x}}\right\rangle\geqslant 0$. Finally, $\delta_{x}$ is not the zero element of $b\mathsf{X}^{\prime}$ because we can always take a $w=\mathbf{1}\in b\mathsf{X}$ with $\left\langle{w,\delta_{x}}\right\rangle=w(x)=1\not=0$. ∎
@@ 2.2. Markov Decision Process
Let $\mathsf{X}$ and $\mathsf{A}$ be metric spaces, as in Section 2.1. A Markov decision process (MDP) with state space $\mathsf{X}$ and action space $\mathsf{A}$ is a tuple $(r,\Gamma,\beta,P)$, where $r$ is a reward function, $x\mapsto\Gamma(x)\subset\mathsf{A}$ is a feasible correspondence, $\beta$ is a discount factor and $P(x,a,\mathop{}\!\mathrm{d}x^{\prime})$ is a distribution over next period states given current state $x$ and action $a$. Let $\mathsf{G}\coloneq\{(x,a)\in\mathsf{X}\times\mathsf{A}:a\in\Gamma(x)\}$. We consider a relatively standard environment, as considered in, say, Bäuerle and Rieder (2011) and Hernández-Lerma and Lasserre (2012), where
@@
    the map $(x,a)\mapsto\int v(x^{\prime})P(x,a,dx^{\prime})$ is continuous on $\mathsf{G}$ whenever $v\in bc\mathsf{X}$.
@@
$v_{\sigma}=\sum_{t=0}^{\infty}(\beta P_{\sigma})^{t}r_{\sigma}=(I-\beta P_{\sigma})^{-1}r_{\sigma}.$ (3)
(See, e.g., Puterman (2014), Theorem 6.1.1.) The lifetime value $v_{\sigma}$ defined in (3) is the unique fixed point in $b\mathsf{X}$ of the policy operator $T_{\sigma}$ defined by $T_{\sigma}\,v=r_{\sigma}+\beta P_{\sigma}\,v$. This operator can be written more explicitly as
@@
(The lifetime value $v_{\sigma}$ is the unique fixed point of $T_{\sigma}$ because the spectral radius of the linear operator $\beta P_{\sigma}$ is $\beta$, so $\beta<1$ implies that $v=r_{\sigma}+\beta P_{\sigma}\,v$ has the unique solution given by the right-hand side of (3).) Iterating on the definition $T_{\sigma}\,v=r_{\sigma}+\beta P_{\sigma}\,v$, we find that
@@
    the value function $v^{*}$ is the unique fixed point of the Bellman operator in $b\mathsf{X}$
@@
    the value function $v^{*}$ is well-defined and contained in $bc\mathsf{X}$ and
@@
See Hernández-Lerma and Lasserre (2012) or Bäuerle and Rieder (2011). ∎
@@
$\int(v^{*}(x^{\prime})-v_{\sigma}(x^{\prime}))P_{\sigma}^{n}(x,\mathop{}\!\mathrm{d}x^{\prime})=0\qquad\text{for all }n\in\mathbbm{N}.$ (6)
@@
Fix $n\in\mathbbm{N}$. Applying the expression for $T_{\sigma}^{n}w$ from (4) twice, first with $v=v_{\sigma}$ and then with $v=v^{*}$, we get
@@ 3.1. Preliminary Results
If $\sigma$ is a feasible policy, then the statements (E1) and (E2) are equivalent.
@@
To show (E1) implies (E2), assume (E1) and fix $x\in\mathsf{X}$ with $v_{\sigma}(x)\geqslant v_{s}(x)$ for all $s\in\Sigma$. Let $\delta_{x}$ be the point evaluation functional generated by $x$. Since $\delta_{x}\in\mathcal{D}(\mathsf{X})$ and $\left\langle{v_{\sigma},\delta_{x}}\right\rangle=v_{\sigma}(x)\geqslant v_{s}(x)=\left\langle{v_{s},\delta_{x}}\right\rangle$ for all $s\in\Sigma$, so (E2) holds. To show (E2) implies (E1), fix $\rho\in\mathcal{D}(\mathsf{X})$ with $\left\langle{v_{\sigma},\rho}\right\rangle\geqslant\left\langle{v_{s},\rho}\right\rangle$ for all $s\in\Sigma$. Suppose to the contrary that for each $x\in\mathsf{X},$ we can find a $\tau\in\Sigma$ such that $v_{\tau}(x)>v_{\sigma}(x)$. Since $v^{*}(x)\geqslant v_{s}(x)$ for all $s\in\Sigma$ and $x\in\mathsf{X}$, we have $v^{*}(x)>v_{\sigma}(x)$ for all $x\in\mathsf{X}$. Hence $\left\langle{v^{*},\rho}\right\rangle>\left\langle{v_{\sigma},\rho}\right\rangle$. This contradiction proves (E1). ∎
@@
To illustrate Lemma 3.2, suppose that $\mathsf{X}$ is discrete. In this case, the lemma tells us that $\left\langle{v_{\sigma},\rho}\right\rangle=\left\langle{v^{*},\rho}\right\rangle$ implies $v_{\sigma}=v^{*}$ (i.e., (E3) holds) when $\rho$ is supported on all of $\mathsf{X}$. In other words, maximizing the scalar performance measure $\left\langle{v_{\sigma},\rho}\right\rangle$ with a highly exploratory initial distribution $\rho$ guarantees global optimality.
The disadvantage of this approach is that evaluating the scalar measure from a highly exploratory initial distribution is costly. For this reason, our main interest is in providing conditions under which (E1) implies (E3). Our conditions relate to irreducibility. The first condition, discussed in Section 3.2, uses a notion of irreducibility from the literature on Banach lattices (see, e.g., Zaanen (2012)), while the second, discussed in Section 4.3, uses an irreducibility concept from the Markov process literature.

### v3

@@ Abstract.
In the theory of dynamic programming, an optimal policy is a policy whose lifetime value dominates that of all other policies from every possible initial condition in the state space. This raises a natural question: when does optimality from a single state imply optimality from every state? Working in a general setting, we provide sufficient conditions for this property that relate to reachability and irreducibility. Our results have significant implications for modern policy-based algorithms used to solve large-scale dynamic programs. We illustrate our findings by applying them to an optimal savings problem via an algorithm that implements gradient ascent in a policy space constructed from neural networks.
@@ 1. Introduction
Dynamic programming (DP) is a major branch of optimization theory with a vast range of applications. Within economics alone, applications extend from monetary and fiscal policy to asset pricing, unemployment, firm investment, wealth dynamics, commodity pricing, sovereign default, the division of labor, natural resource extraction, human capital accumulation, retirement decisions, portfolio choice, and dynamic pricing. Dynamic programs that include uncertainty are often called Markov decision processes (MDPs) and the theory of such processes has been extensively developed (see, e.g., Bäuerle and Rieder (2011), Hernández-Lerma and Lasserre (2012), Bertsekas (2012), or Bertsekas (2022)).
Much of the recent surge in interest in MDPs has been fueled by artificial intelligence and reinforcement learning (see, e.g., Bertsekas (2021) or Kochenderfer et al. (2022)). In recent years, researchers solving large-scale MDPs have moved away from value-based methods and towards policy-based methods, such as policy gradient ascent (see, e.g., Sutton et al. (1999), Lan et al. (2023), Kumar et al. (2023), Murphy (2024)). These policy-based methods seek to maximize a real-valued objective, such as
@@
Despite these successes, there is one significant disadvantage of policy gradient methods: unlike traditional DP algorithms, these methods are not guaranteed to converge to an optimal policy, even in relatively straightforward settings. To understand the issues at hand, recall that an optimal policy is a $\sigma\in\Sigma$ such that $v_{\sigma}(x)=\max_{s\in\Sigma}v_{s}(x)$ for every $x\in\mathsf{X}$; that is, a policy $\sigma$ such that following this policy in every period leads to maximum lifetime value from every initial state $x$. Even if one attains a global maximum in (1), with maximization over all $\sigma\in\Sigma$, there is no guarantee that the resulting policy will be an optimal policy. Indeed, the result depends in general on the initial distribution $\rho$, which is not one of the primitives of the DP problem.
Regarding this issue, the prevailing view is that relative optimality depends on the initial distribution $\rho$ having large support. For example, in a major study of policy gradient methods, Bhandari and Russo (2024) state that “Policy gradient methods have poor convergence properties if applied without an exploratory initial distribution” (Bhandari and Russo, 2024, p. 1910).
One disadvantage of choosing $\rho$ with large support is that computing the integral (1) becomes computationally expensive, since the function $v_{\sigma}(x)$ is expensive to evaluate. Another is that, in general, this integral cannot be computed exactly, and, moreover, the implications of a relatively poor approximation to the integral for optimality theory are not well-known. A third issue is that existing conditions for “sufficiently large supports” are difficult to test in practice, making it hard to understand whether or not the relevant theory applies. (We return to this last point below.)
These observations motivate the following question: Does there exist a large and important class of models for which it is permissible to choose $\rho$ with small support, while still maintaining the property that maximization of this criterion over all $\sigma\in\Sigma$ yields a (globally) optimal policy? At the extreme, when can $\rho$ be concentrated at a single point $x$, so that the maximization criterion $m(\sigma)$ from (1) is just $v_{\sigma}(x)$? In other words, when does optimality at a single state imply optimality at every state?
We show that, for standard MDPs on general state spaces, strong irreducibility of the Markov dynamics generated by $\sigma$ is sufficient for this property. Specifically, if a policy is optimal at a single state and has a strongly irreducible transition kernel, then this optimality propagates throughout the entire state space, making the policy globally optimal. (Here “strong irreducibility” refers to irreducibility of positive operators under the standard Riesz space definition.) In addition, we show that weaker forms of irreducibility are also sufficient for this result when $\sigma$ is also continuous.
@@
Papers examining theoretical properties of policy gradient methods have some connection to our work. For example, Bhandari and Russo (2024) examine when policy gradient ascent (actually decent) yields a globally optimal policy. To step from local to global optimality, they restrict the classes of MDPs under consideration and require that the initial distribution $\rho$ dominates the discounted state occupancy measure under the optimal policy (i.e, the measure is absolutely continuous with respect to $\rho$). This condition is difficult to verify in practice, since the optimal policy is a priori unknown (and, in fact, the object of the whole computational procedure). Here we avoid both large support restrictions on $\rho$ and explicit restrictions on the optimal policy. (At the same time, Bhandari and Russo (2024) supply many valuable new results concerning the policy gradient algorithm, and we make no direct contribution to this analysis.)
Other papers that examine theoretical properties of gradient policy methods include Khodadadian et al. (2021), Agarwal et al. (2021), and Xiao (2022). However, in these papers, the focus is on proving the convergence of $\int v_{\sigma}\mathop{}\!\mathrm{d}\rho$ to its maximal value, rather than obtaining global convergence from local convergence (as we do here). At the same time, these papers provide valuable rates of convergence for specific algorithms, which we do not discuss. A related line of research focuses on average-optimal policies in finite-state MDPs by leveraging specific state space structures under some policies. This includes exploring unichain, multichain, communicating, and weakly communicating MDPs to study algorithmic convergence (Bartlett and Tewari, 2009; Puterman, 2014). Our results focus on the continuous state setting, where gradient policy methods are usually applied.
In an application connected to the income fluctuation problem, we examine policy gradient methods under different irreducibility settings using neural network representations of policy functions. In this we follow a recent trend in economics towards solving large scale dynamic problems using neural networks to represent equilibrium objects (see, e.g., Maliar et al. (2021) or Friedl et al. (2023)). However, these papers do not examine the link between local and global optimality. In fact they do not use the class of policy gradient methods described above, focusing instead on minimizing Euler or Bellman residuals.
The paper is structured as follows. Section 2 provides background on MDPs. Section 3 states our main results in a general setting. Section 4 examines how the irreducibility condition from Section 3 can be weakened while still obtaining some transmission of optimality across states. Section 5.1 illustrates our theoretical results in the context of a benchmark optimal savings problem. Section 6 outlines avenues for future work.
@@ 2.1. Preliminaries
A transition kernel on $\mathsf{X}$ is a function $Q$ from $\mathsf{X}\times\mathscr{B}$ to $[0,1]$ such that $x\mapsto Q(x,B)$ is Borel measurable for all $B\in\mathscr{B}$ and $B\mapsto Q(x,B)$ is a Borel probability measure for all $x\in\mathsf{X}$. To any such kernel $Q$ we associate a bounded linear operator on $b\mathsf{X}$, often referred to as its Markov operator and also denoted by $Q$, via
$f\mapsto Qf,\qquad(Qf)(x)=\int f(x^{\prime})Q(x,\mathop{}\!\mathrm{d}x^{\prime}).$ (2)
Typically, $(Qf)(x)$ represents the expectation of $f(X_{t+1})$ given that $X_{t}=x$ and $X_{t+1}$ is drawn from $Q(x,\mathop{}\!\mathrm{d}x^{\prime})$.
As usual, the positive cone of $b\mathsf{X}$, denoted here by $b\mathsf{X}_{+}$, is all $v\in b\mathsf{X}$ with $v\geqslant 0$. Let $b\mathsf{X}^{\prime}$ be the dual space of $b\mathsf{X}$ and let $b\mathsf{X}^{\prime}_{+}$ to be the positive cone of $b\mathsf{X}^{\prime}$. The set $b\mathsf{X}^{\prime}_{+}$ contains, among other objects, the set $\mathcal{D}(\mathsf{X})$ of Borel probability measures on $\mathsf{X}$. For simplicity, elements of $\mathcal{D}(\mathsf{X})$ are referred to as distributions. For $\rho\in\mathcal{D}(\mathsf{X})$ and $f\in b\mathsf{X}$ we set
@@
For each $x\in\mathsf{X}$, the point evaluation functional generated by $x$ is the map $\delta_{x}$ that sends $w\in b\mathsf{X}$ into $w(x)\in\mathbbm{R}$. Below it will be convenient for us to write this in dual notation, so that $\left\langle{w,\delta_{x}}\right\rangle=w(x)$ for every $w\in b\mathsf{X}$. We will make use of the following lemma.
@@
Regarding continuity, if $w_{n}\to w$ in $b\mathsf{X}$, then $w_{n}\to w$ pointwise on $\mathsf{X}$, so $\left\langle{w_{n},\delta_{x}}\right\rangle=w_{n}(x)\to w(x)=\left\langle{w,\delta_{x}}\right\rangle$. Regarding positivity, it suffices to show that $\left\langle{w,\delta_{x}}\right\rangle\geqslant 0$ whenever $w\geqslant 0$. This clearly holds, since $w\geqslant 0$ implies $w(x)=\left\langle{w,\delta_{x}}\right\rangle\geqslant 0$. Finally, $\delta_{x}$ is not the zero element of $b\mathsf{X}^{\prime}$ because $w=\mathbf{1}$ is in $b\mathsf{X}$ and $\left\langle{w,\delta_{x}}\right\rangle=w(x)=1\not=0$. ∎
@@ 2.2. Markov Decision Process
Let $\mathsf{X}$ and $\mathsf{A}$ be metric spaces, as in Section 2.1. A Markov decision process (MDP) with state space $\mathsf{X}$ and action space $\mathsf{A}$ is a tuple $(r,\Gamma,\beta,P)$, where $r$ is a reward function, $x\mapsto\Gamma(x)\subset\mathsf{A}$ is a feasible correspondence, $\beta\in\mathbbm{R}$ is a discount factor and $P(x,a,\mathop{}\!\mathrm{d}x^{\prime})$ is a distribution over next period states given current state $x$ and action $a$. Let $\mathsf{G}\coloneq\{(x,a)\in\mathsf{X}\times\mathsf{A}:a\in\Gamma(x)\}$. We consider a relatively standard environment, as considered in, say, Bäuerle and Rieder (2011) and Hernández-Lerma and Lasserre (2012), where
@@
    the map $(x,a)\mapsto\int v(x^{\prime})P(x,a,dx^{\prime})$ is continuous on $\mathsf{G}$ whenever $v\in cb\mathsf{X}$.
@@
$v_{\sigma}=\sum_{t=0}^{\infty}(\beta P_{\sigma})^{t}r_{\sigma}$ (3)
(see, e.g., Puterman (2014), Theorem 6.1.1). We will use the fact that $v_{\sigma}$ is also the unique fixed point in $b\mathsf{X}$ of the policy operator $T_{\sigma}$ defined by $T_{\sigma}\,v=r_{\sigma}+\beta P_{\sigma}\,v$. This operator can be written more explicitly as
@@
(The lifetime value $v_{\sigma}$ is the unique fixed point of $T_{\sigma}$ because the spectral radius of the linear operator $\beta P_{\sigma}$ is $\beta$, so, by the geometric series lemma, $\beta<1$ implies that $v=r_{\sigma}+\beta P_{\sigma}\,v$ has the unique solution given by the right-hand side of (3).) Iterating on the definition $T_{\sigma}\,v=r_{\sigma}+\beta P_{\sigma}\,v$, we find that
@@
    the value function $v^{*}$ is the unique fixed point of the Bellman operator in $b\mathsf{X}$,
@@
    the value function $v^{*}$ is well-defined and contained in $cb\mathsf{X}$, and
@@
See Bäuerle and Rieder (2011) or Hernández-Lerma and Lasserre (2012). ∎
@@
$\sum_{n\in\mathbbm{N}}\int(v^{*}(x^{\prime})-v_{\sigma}(x^{\prime}))P_{\sigma}^{n}(x,\mathop{}\!\mathrm{d}x^{\prime})=0.$ (6)
@@
Fix $n\in\mathbbm{N}$. Applying the expression for $T_{\sigma}^{n}v$ from (4) twice, first with $v=v_{\sigma}$ and then with $v=v^{*}$, we get
@@ 3.1. Preliminary Results
For all $\sigma\in\Sigma$, the statements (E1) and (E2) are equivalent.
@@
To show (E1) implies (E2), assume (E1) and fix $x\in\mathsf{X}$ with $v_{\sigma}(x)\geqslant v_{s}(x)$ for all $s\in\Sigma$. Let $\delta_{x}$ be the point evaluation functional generated by $x$. Since $\delta_{x}\in\mathcal{D}(\mathsf{X})$ and $\left\langle{v_{\sigma},\delta_{x}}\right\rangle=v_{\sigma}(x)\geqslant v_{s}(x)=\left\langle{v_{s},\delta_{x}}\right\rangle$ for all $s\in\Sigma$, so (E2) holds. To show (E2) implies (E1), fix $\rho\in\mathcal{D}(\mathsf{X})$ with $\left\langle{v_{\sigma},\rho}\right\rangle\geqslant\left\langle{v_{s},\rho}\right\rangle$ for all $s\in\Sigma$. Since at least one $s\in\Sigma$ is optimal (Proposition 2.2), this yields $\left\langle{v_{\sigma},\rho}\right\rangle\geqslant\left\langle{v^{*},\rho}\right\rangle$. Now suppose to the contrary that for each $x\in\mathsf{X},$ we can find a $\tau\in\Sigma$ such that $v_{\tau}(x)>v_{\sigma}(x)$. Since $v^{*}(x)\geqslant v_{s}(x)$ for all $s\in\Sigma$ and $x\in\mathsf{X}$, we have $v^{*}(x)>v_{\sigma}(x)$ for all $x\in\mathsf{X}$. Hence $\left\langle{v^{*},\rho}\right\rangle>\left\langle{v_{\sigma},\rho}\right\rangle$. This contradiction proves (E1). ∎
@@
Loosely speaking, this lemma shows that maximizing the scalar performance measure $\left\langle{v_{\sigma},\rho}\right\rangle$ with a highly exploratory initial distribution $\rho$ promotes global optimality.
The disadvantage of this approach is that evaluating the scalar measure from a such an initial distribution is costly. For this reason, our main interest is in providing conditions under which (E1) implies (E3). Our conditions relate to irreducibility. The first condition, discussed in Section 3.2, uses a notion of irreducibility from the literature on Banach lattices (see, e.g., Zaanen (2012)), while the second, discussed in Section 4.3, uses an irreducibility concept from the Markov process literature.